OpenStudy (cookiemonster18):

1.The pressure, P of a gas varies inversely with its volume, V. Pressure is measured in units of Pa. Suppose that particular amount of gas is initially at a pressure of 84 Pa at a volume of 36 L. If volume is expanded is expanded to 216, what will the new pressure be? A.6 Pa B.14 Pa C.216 Pa*** D.504 Pa

OpenStudy (wolf1728):

This is Boyle's Law P1*V1 = P2*V2 84 * 36 = P2 * 216 P2 = (84 * 36) / 216 Can you solve it from there?

OpenStudy (cookiemonster18):

Do the data in the table represent a direct variation or an inverse variation? Write an equation to model the data in the table. X ; 1 , 3 , 5 , 10 Y ; 4 , 12 , 20 , 40 A.direct variation; y=4x B.direct variation; xy=1/4 C.inverse variation; xy=4 D.inverse variation; xy=1/4

OpenStudy (cookiemonster18):

@wolf1728 not really, I haven't learned how to solve these type of problems but I'll give it a try can you help with problem #2 as I do

OpenStudy (wolf1728):

Maybe this is helping a bit too much, but here is a Boyle's Law Calculator: http://www.1728.org/boyle.htm

OpenStudy (cookiemonster18):

#1 is B 14Pa am I correct @wolf1728

OpenStudy (wolf1728):

Yes it is 14 Pascals

OpenStudy (cookiemonster18):

Can you please help me with #2. @wolf1738 its my last question

OpenStudy (wolf1728):

Sure it is pretty easy For each value of y , check it against the corresponding value of X - see anything they have in common?